OpenStudy (anonymous):

G(x) = x^2 + 6x + 1 Find the solutions of g(x). Show each step

OpenStudy (anonymous):

g(x) = x2 +6x + 1 -1 -1 -1 = x^2 + 6x 6/2 = 3 ^2 = 9 +9 +9 8 = (x+3) ^2 -8 -8 It's Here that i don't understand the square root part.

OpenStudy (anonymous):

@MDoodler ?

OpenStudy (anonymous):

Use the quadratic formula on x^2 + 6x + 1 \(x = \frac{ - b \sqrt {b^2 - 4ac}}{2a} \)

OpenStudy (anonymous):

\(x = \frac{ -6 \sqrt {6^2 - 4a(1)(1)}}{2a} \)

OpenStudy (anonymous):

:/ huh?

OpenStudy (anonymous):

Ok, you have not used the quadratic formula yet in class?

OpenStudy (anonymous):

I never understood it so i use another way. the one above.

OpenStudy (anonymous):

Your solutions are going to be roots so it is best to use the quadratic formula

OpenStudy (anonymous):

square roots that is

OpenStudy (anonymous):

could you show me how to use it?

OpenStudy (anonymous):

Yes, sorry my computer went out.

OpenStudy (anonymous):

From here 8 = (x+3)^2, which you already have, we need to square both sides \( 8 = (x+3) ^2\) Square both sides \( \sqrt{ 8} = \sqrt{(x+3) ^2}\) \( \sqrt{ 8} = x+3\) \( \sqrt{ 8} \) \( \sqrt{ 8}\) We squared the right side now lets square the left side To square the left side, we need to simplify the square \( \sqrt{ 8}\) \( \sqrt{ 4} \sqrt{2} \) \( 2 \sqrt{2} \) Now we have \( 2 \sqrt{2} = x+ 3\) \( -3 + 2 \sqrt{2} = x+ 3 + -3\) \( -3 + 2 \sqrt{2} = x\) So your solutions are \( x = -3 + 2 \sqrt{2} \) \( x = 2 \sqrt{2} -3 \) I can also show you the quadratic formula way too, which I think is easier

OpenStudy (anonymous):

Sure, can you show the easier way. But I do understand this method.

OpenStudy (anonymous):

Thank you so much ^_^